OpenStudy (anonymous):

solve the equation: 42=-3(m+6)+m

OpenStudy (anonymous):

Answer is -30. Need steps.

OpenStudy (anonymous):

42 = -3(m + 6) + m 42 = -3m - 18 + m 42 = -2m - 18 42 + 18 = -2m - 18 + 18 60 = -2m 60/-2 = -2m/-2 m = -30 hope that helped :)

OpenStudy (anonymous):

Thank You so much :)

OpenStudy (anonymous):

anytime :)